# Plan Then Retrieve: Reinforcement Learning-Guided Complex Reasoning over Knowledge Graphs

### v2

@@ Abstract.
Relevance for WWW. Graph-RFT is a key task for the integration of LLMs, KGs and web application, which is a scientific challenge within the scope of the Semantics and Knowledge Track.
@@ 1. Introduction
Knowledge Graph Question Answering (KGQA) aims to answer natural language questions by reasoning over structured knowledge graphs (KGs) that store explicit, factual knowledge in the form of triples. KGQA plays a central role in web-based intelligent systems such as search, recommendation, and social platforms (Khorashadizadeh et al., 2024; Zhao et al., 2024; Zeng et al., 2024), where accurate and explainable reasoning over structured data is crucial. Recent advances in large language models (LLMs) have transformed natural language understanding, reasoning, and generation (Hendrycks et al., 2020; Clark et al., 2018; Guo et al., 2025; Shi et al., 2025; Liu et al., 2025a). Leveraging their strong generalization and reasoning capabilities, recent KGQA methods increasingly rely on LLMs, integrating knowledge graphs as structured, verifiable knowledge sources to enhance factual precision and interpretability (Sun et al., 2023; Jiang et al., 2023; Tan et al., 2025; Ma et al., 2025; Liu et al., 2025b; Xue et al., 2024).
Existing LLM-based KGQA approaches fall into two paradigms: semantic parsing (SP), which translates natural questions into executable logical queries (e.g., SPARQL) and runs them over a KG (Li et al., 2023a; Luo et al., 2023b; Xu et al., 2024), and retrieval-augmented generation (RAG), which retrieves relevant KG triples to condition an LLM’s generation (Tan et al., 2025; Xu et al., 2024). Despite substantial progress, both paradigms struggle in complex settings for two main reasons. First, they lack a reliable mechanism to judge whether a KG contains the facts necessary to answer a question. SP and RAG methods commonly assume that all required triples are present in the KG, a strong and unrealistic assumption for manually curated graphs. When coverage is incomplete, SP fails to produce correct executable queries, and RAG offers no built-in way to refine retrieval or to decide when to consult external (unstructured) sources. Second, these methods often lack the ability to perform coherent multi-step planning and adaptive retrieval scheduling. For example in Figure 1(2), an LLM may correctly decompose a question into “Which screenwriters wrote Analyze That?” and retrieve “Harold Ramis,” but then fail to continue the planned decomposition and instead reformulates the follow-up as “When was Harold Ramis released?”, confusing the person with a film title. This failure stems from constrained local decision-making and an absence of coherent multi-step planning, which leads to retrieval and reasoning errors even when the KG contains the necessary facts. The comparison for previous work and Graph-RFT is in Table 1.
@@ 2. Related Work
SP Methods convert natural language questions into structural queries (e.g., SPARQL) that can be executed on KGs to retrieve answers. Early research efforts (Hu et al., 2017; Xiong et al., 2017; Luo et al., 2025b) in semantic parsing often involved generating query graphs, resembling subgraphs of the KG and allow for direct mapping to logical forms. Lan (Lan and Jiang, 2020) incorporates constraints into the query graph, which effectively narrows the search space and facilitates more flexible query graph generation. Subsequent works (Atif et al., 2023; Schick et al., 2023; Raffel et al., 2020; Wei et al., 2022; Jiang et al., 2022; Lan and Jiang, 2020) utilize sequence-to-sequence models for multi-hop path and SPARQL-expression generation, enhancing the semantic parsing process. Recently, with the emergence of LLMs, various methods unify KGs and LLMs for KGQA.ChatKBQA (Luo et al., 2023b) generates question-aligned logical forms by treating LLMs as semantic parsers and retrieves entities and relations from KGs. RGR-KBQA (Feng and He, 2025) retrieves factual knowledge from KGs to enhance the semantic understanding capabilities of LLMs and finetune them to generate the logical form. Rule-KBQA (Zhang et al., 2025) guides logical form generation via a two-phase process that inducts a rule library and deducts forms via a rule-guided agent. However, these methods depend heavily on the assumption that KGs are complete, which means no answers can be obtained if information is missing.
@@
RA Methods retrieve relevant KG triple based on input questions, then use these triples to guide LLMs to generate accurate answers, emphasizing the dynamic interaction between KG retrieval and LLM reasoning. ToG (Sun et al., 2023) employs an explore-and-exploit approach, which can discover the most promising reasoning paths and return the most likely reasoning results. GoG (Xu et al., 2024) proposes a think-search-generate paradigm, alleviating the incompleteness issue of KGs by using LLMs’ internal knowledge. PoG (Tan et al., 2025) leverages knowledge reasoning paths as retrieval-augmented inputs for LLMs, thereby alleviating issues related to insufficient domain knowledge and unfaithful reasoning chains. DoG (Ma et al., 2025) introduces a subgraph-focused mechanism that allows LLMs to attempt intermediate answers after each reasoning step, effectively mitigating the drawbacks of lengthy reasoning trajectories. However, most of these approaches depend on powerful closed-source LLM APIs (e.g., GPT-4 (Achiam et al., 2023)), which results in substantial performance degradation when weaker open-source models are used as backbones. KG-Agent (Jiang et al., 2024) develops an iterative reasoning framework that integrates an LLM, a multifunctional toolbox, a KG-based executor, and a knowledge memory module, enabling smaller LLMs to make decisions autonomously. SymAgent (Liu et al., 2025b) conceptualizes the KG as a dynamic environment and is designed to autonomously and effectively integrate the capabilities of both the LLM and the KG. Graph-R1 (Luo et al., 2025a) proposes lightweight knowledge hypergraph construction and retrieval framed as a multi-turn agent–environment interaction, optimizing the reasoning process via an end-to-end reward mechanism. DynaSearcher (Hao et al., 2025) utilizes knowledge graphs as external structured knowledge to guide the search process by explicitly modeling entity relationships, thereby ensuring factual consistency in intermediate queries.
@@ 3.2. CoT Fine-Tuning for Reasoning Activation
where is the question, is the concatenated sequence of reasoning steps and the answer, is the model’s token distribution. The output model serves as the initialization for the next stage, ensuring a robust foundation for RL.
@@ 3.3.1. Rollout with Search Tools
Retrieve Facts from the KG. For each single steps obtained through planning, we need to sequentially select relevant triples from the KG. To address this, we designe two interactive toolkits.
@@
Given the obtained single-hop questions (e.g., “Identify the country that uses the Iranian Rail”), we first invoke the Relation Search Tool to retrieve a candidate set of relations associated with the initial (topic) entity. The LLM then reasons over this set and selects the most appropriate relation, for example, choosing “used” instead of “used_in” in the retrieval function. This greedy strategy simplifies the reasoning process by selecting a locally optimal relation at each step, thereby reducing reliance on global contextual information. Next, the Neighbor Search Tool is called to obtain specific triples and their corresponding answers (e.g., “The Islamic Republic of Iran”). If the LLM determines that the retrieved triples are insufficient to answer the question, or if no answer is found in the knowledge graph (i.e., when ¡neighbor_information¿ and ¡/neighbor_information¿ return “No information in the KG, please use web tool”), the system directly invokes the search engine. The answers to higher-priority (according to the order produced by the subquestion plan) single-hop questions can then be combined with those of lower-priority ones to generate new single-hop questions. By iteratively repeating this process, the system is able to derive answers to complex multi-hop questions. Notably, if the maximum iteration limit is reached without successfully generating an answer, the parametric knowledge of the LLM is used to provide a final response.
@@ 3.3.2. Reward Design
| RoG (Luo et al., 2023a) | SL | Qwen2.5-7B-base | 63.7 | 53.6 | 84.2 | 75.3 | - | - | - | - |
| ChatKBQA (Luo et al., 2023b) | SL | Qwen2.5-7B-base | 72.4 | 37.8 | 75.5 | 47.1 | - | - | - | - |
@@ 4. Experiments
Datasets. We evaluate Graph-RFT on four KGQA datasets, including three multi-hop datasets: CWQ (Talmor and Berant, 2018), WebQSP (Yih et al., 2016), GrailQA (Gu et al., 2021) and one single-hop dataset: SimpleQuestions (Petrochuk and Zettlemoyer, [n. d.]). We consider two scenarios: Complete Knowledge Graph (CKG) and Incomplete Knowledge Graph (IKG). For CKG, we directly use the datasets provided by GoG (Xu et al., 2024). For IKG, since GoG only provides CWQ and WebQSP, we construct IKG versions of GrailQA and SimpleQuestions by sampling 3,000 multi-hop questions from GrailQA and 3,000 questions from SimpleQuestions. We further generate four IKGs with varying completeness levels IKG-20%, IKG-40%, IKG-60%for example, IKG-40% removes 40% of the critical triples for each question and all relations between the corresponding entity pairs. Freebase (Bollacker et al., 2008) serves as the background KG for all datasets. The impact of using alternative knowledge graphs is discussed in Appendix A.5.
@@ 4.1. Main Results (RQ1)
Table 2 presents our main results on four KGQA datasets. We can observe that Graph-RFT consistently achieves SOTA performance across all settings, even when built upon smaller backbones such as Qwen2.5-7B-base, outperforming GPT-4-based methods.
@@ 4.3. Different KG Incompleteness (RQ3)
Impact of Rewards with or without Penalty. From Figure 4, Graph-RFT increases the frequency of external searches as KG completeness decreases. This behavior reflects the model’s ability to proactively compensate for missing knowledge required for reasoning through web retrieval. Moreover, the model’s search strategy varies notably with question complexity: compared to WebQSP (maximum 2-hop), Graph-RFT conducts significantly more external searches when tackling the more complex multi-hop questions in CWQ (maximum 4-hop). This demonstrates that the model can dynamically adjust its web search frequency according to question complexity, effectively aligning knowledge acquisition with the reasoning demands of the task. In addition, Graph-RFT can also save time and improve efficiency.
Impact of Different KG Incompleteness. From Figure 5, we can observe that Graph-RFT consistently outperforms GoG across all degrees of incompleteness. This improvement arises because GoG is primarily confined to local decision-making, lacking the global and hierarchical planning needed for effective question understanding and tool invocation in complex problem decomposition. In contrast, when KG triples are missing, GoG relies solely on the LLM’s internal knowledge to generate answers, whereas our model retrieves relevant external documents through web search, thereby supplementing incomplete KG information and improving answer accuracy. Notably, the improvement on CWQ is much more pronounced than on WebQSP, validating Graph-RFT’s strong planning ability and logical decomposition mechanism for tackling complex reasoning tasks.
@@ Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.References
- (1)
- Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774 (2023).
@@
- Atif et al. (2023) Farah Atif, Ola El Khatib, and Djellel Difallah. 2023. Beamqa: Multi-hop knowledge graph question answering with sequence-to-sequence prediction and beam search. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval. 781–790.
@@
- Baek et al. (2023) Jinheon Baek, Alham Fikri Aji, and Amir Saffari. 2023. Knowledge-augmented language model prompting for zero-shot knowledge graph question answering. arXiv preprint arXiv:2306.04136 (2023).
@@
- Bollacker et al. (2008) Kurt Bollacker, Colin Evans, Praveen Paritosh, Tim Sturge, and Jamie Taylor. 2008. Freebase: a collaboratively created graph database for structuring human knowledge. In Proceedings of the 2008 ACM SIGMOD international conference on Management of data. 1247–1250.
- Clark et al. (2018) Peter Clark, Isaac Cowhey, Oren Etzioni, Tushar Khot, Ashish Sabharwal, Carissa Schoenick, and Oyvind Tafjord. 2018. Think you have solved question answering? try arc, the ai2 reasoning challenge. arXiv preprint arXiv:1803.05457 (2018).
- Descartes (1901) René Descartes. 1901. Discourse on method. Aladdin Book Company.
@@
- Feng and He (2025) Tengfei Feng and Liang He. 2025. Rgr-kbqa: Generating logical forms for question answering using knowledge-graph-enhanced large language model. In Proceedings of the 31st International Conference on Computational Linguistics. 3057–3070.
@@
- Gu et al. (2021) Yu Gu, Sue Kase, Michelle Vanni, Brian Sadler, Percy Liang, Xifeng Yan, and Yu Su. 2021. Beyond iid: three levels of generalization for question answering on knowledge bases. In Proceedings of the web conference 2021. 3477–3488.
- Guo et al. (2025) Daya Guo, Dejian Yang, Haowei Zhang, Junxiao Song, Ruoyu Zhang, Runxin Xu, Qihao Zhu, Shirong Ma, Peiyi Wang, Xiao Bi, et al. 2025. Deepseek-r1: Incentivizing reasoning capability in llms via reinforcement learning. arXiv preprint arXiv:2501.12948 (2025).
@@
- Hao et al. (2025) Chuzhan Hao, Wenfeng Feng, Yuewei Zhang, and Hao Wang. 2025. DynaSearcher: Dynamic Knowledge Graph Augmented Search Agent via Multi-Reward Reinforcement Learning. arXiv preprint arXiv:2507.17365 (2025).
- Hendrycks et al. (2020) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2020. Measuring massive multitask language understanding. arXiv preprint arXiv:2009.03300 (2020).
@@
- Hu et al. (2017) Sen Hu, Lei Zou, Jeffrey Xu Yu, Haixun Wang, and Dongyan Zhao. 2017. Answering natural language questions by subgraph matching over knowledge graphs. IEEE Transactions on Knowledge and Data Engineering 30, 5 (2017), 824–837.
@@
- Jiang et al. (2023) Jinhao Jiang, Kun Zhou, Zican Dong, Keming Ye, Wayne Xin Zhao, and Ji-Rong Wen. 2023. Structgpt: A general framework for large language model to reason over structured data. arXiv preprint arXiv:2305.09645 (2023).
@@
- Jiang et al. (2024) Jinhao Jiang, Kun Zhou, Wayne Xin Zhao, Yang Song, Chen Zhu, Hengshu Zhu, and Ji-Rong Wen. 2024. Kg-agent: An efficient autonomous agent framework for complex reasoning over knowledge graph. arXiv preprint arXiv:2402.11163 (2024).
@@
- Jiang et al. (2022) Jinhao Jiang, Kun Zhou, Wayne Xin Zhao, and Ji-Rong Wen. 2022. Unikgqa: Unified retrieval and reasoning for solving multi-hop question answering over knowledge graph. arXiv preprint arXiv:2212.00959 (2022).
@@
- Jin et al. (2025) Bowen Jin, Hansi Zeng, Zhenrui Yue, Jinsung Yoon, Sercan Arik, Dong Wang, Hamed Zamani, and Jiawei Han. 2025. Search-r1: Training llms to reason and leverage search engines with reinforcement learning. arXiv preprint arXiv:2503.09516 (2025).
- Khorashadizadeh et al. (2024) Hanieh Khorashadizadeh, Fatima Zahra Amara, Morteza Ezzabady, Frédéric Ieng, Sanju Tiwari, Nandana Mihindukulasooriya, Jinghua Groppe, Soror Sahri, Farah Benamara, and Sven Groppe. 2024. Research trends for the interplay between large language models and knowledge graphs. arXiv preprint arXiv:2406.08223 (2024).
@@
- Lan and Jiang (2020) Yunshi Lan and Jing Jiang. 2020. Query graph generation for answering multi-hop complex questions from knowledge bases. Association for Computational Linguistics.
@@
- Li et al. (2023a) Tianle Li, Xueguang Ma, Alex Zhuang, Yu Gu, Yu Su, and Wenhu Chen. 2023a. Few-shot in-context learning for knowledge base question answering. arXiv preprint arXiv:2305.01750 (2023).
@@
- Li et al. (2023b) Xingxuan Li, Ruochen Zhao, Yew Ken Chia, Bosheng Ding, Lidong Bing, Shafiq Joty, and Soujanya Poria. 2023b. Chain of knowledge: A framework for grounding large language models with structured knowledge bases. arXiv preprint arXiv:2305.13269 3 (2023).
@@
- Liu et al. (2025b) Ben Liu, Jihai Zhang, Fangquan Lin, Cheng Yang, Min Peng, and Wotao Yin. 2025b. Symagent: A neural-symbolic self-learning agent framework for complex reasoning over knowledge graphs. In Proceedings of the ACM on Web Conference 2025. 98–108.
- Liu et al. (2025a) Ziyu Liu, Zeyi Sun, Yuhang Zang, Xiaoyi Dong, Yuhang Cao, Haodong Duan, Dahua Lin, and Jiaqi Wang. 2025a. Visual-rft: Visual reinforcement fine-tuning. arXiv preprint arXiv:2503.01785 (2025).
- Luo et al. (2025a) Haoran Luo, Guanting Chen, Qika Lin, Yikai Guo, Fangzhi Xu, Zemin Kuang, Meina Song, Xiaobao Wu, Yifan Zhu, Luu Anh Tuan, et al. 2025a. Graph-r1: Towards agentic graphrag framework via end-to-end reinforcement learning. arXiv preprint arXiv:2507.21892 (2025).
@@
- Luo et al. (2023b) Haoran Luo, Zichen Tang, Shiyao Peng, Yikai Guo, Wentai Zhang, Chenghao Ma, Guanting Dong, Meina Song, Wei Lin, Yifan Zhu, et al. 2023b. Chatkbqa: A generate-then-retrieve framework for knowledge base question answering with fine-tuned large language models. arXiv preprint arXiv:2310.08975 (2023).
@@
- Luo et al. (2025b) Linhao Luo, Jiaxin Ju, Bo Xiong, Yuan-Fang Li, Gholamreza Haffari, and Shirui Pan. 2025b. Chatrule: Mining logical rules with large language models for knowledge graph reasoning. In Pacific-Asia Conference on Knowledge Discovery and Data Mining. Springer, 314–325.
@@
- Luo et al. (2023a) Linhao Luo, Yuan-Fang Li, Gholamreza Haffari, and Shirui Pan. 2023a. Reasoning on graphs: Faithful and interpretable large language model reasoning. arXiv preprint arXiv:2310.01061 (2023).
@@
- Ma et al. (2025) Jie Ma, Zhitao Gao, Qi Chai, Wangchun Sun, Pinghui Wang, Hongbin Pei, Jing Tao, Lingyun Song, Jun Liu, Chen Zhang, et al. 2025. Debate on graph: a flexible and reliable reasoning framework for large language models. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 39. 24768–24776.
@@
- Petrochuk and Zettlemoyer ([n. d.]) M Petrochuk and L Zettlemoyer. [n. d.]. Simplequestions nearly solved: A new upperbound and baseline approach. arXiv 2018. arXiv preprint arXiv:1804.08798 ([n. d.]).
@@
- Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. 2020. Exploring the limits of transfer learning with a unified text-to-text transformer. Journal of machine learning research 21, 140 (2020), 1–67.
- Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Eric Hambro, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language models can teach themselves to use tools. Advances in Neural Information Processing Systems 36 (2023), 68539–68551.

### v3

@@ 1. Introduction
Knowledge Graph Question Answering (KGQA) aims to answer natural language questions by reasoning over structured knowledge graphs (KGs) that store explicit, factual knowledge in the form of triples. KGQA plays a central role in web-based intelligent systems such as search, recommendation, and social platforms (Khorashadizadeh et al., 2024; Zhao et al., 2024; Zeng et al., 2024; Xie et al., 2023, 2024, 2025; Wang et al., 2023), where accurate and explainable reasoning over structured data is crucial. Recent advances in large language models (LLMs) have transformed natural language understanding, reasoning, and generation (Hendrycks et al., 2020; Clark et al., 2018; Guo et al., 2025; Shi et al., 2025; Liu et al., 2025b). Leveraging their strong generalization and reasoning capabilities, recent KGQA methods increasingly rely on LLMs, integrating knowledge graphs as structured, verifiable knowledge sources to enhance factual precision and interpretability (Sun et al., 2023; Jiang et al., 2023; Tan et al., 2025b; Ma et al., 2025; Liu et al., 2025a; Xue et al., 2024).
Existing LLM-based KGQA approaches fall into two paradigms: semantic parsing (SP), which translates natural questions into executable logical queries (e.g., SPARQL) and runs them over a KG (Li et al., 2023a; Luo et al., 2023a; Xu et al., 2024), and retrieval-augmented generation (RAG), which retrieves relevant KG triples to condition an LLM’s generation (Tan et al., 2025b; Xu et al., 2024). Despite substantial progress, both paradigms struggle in complex settings for two main reasons. First, they lack a reliable mechanism to judge whether a KG contains the facts necessary to answer a question. SP and RAG methods commonly assume that all required triples are present in the KG, a strong and unrealistic assumption for manually curated graphs. When coverage is incomplete, SP fails to produce correct executable queries, and RAG offers no built-in way to refine retrieval or to decide when to consult external (unstructured) sources. Second, these methods often lack the ability to perform coherent multi-step planning and adaptive retrieval scheduling. For example in Figure 1(2), an LLM may correctly decompose a question into “Which screenwriters wrote Analyze That?” and retrieve “Harold Ramis,” but then fail to continue the planned decomposition and instead reformulates the follow-up as “When was Harold Ramis released?”, confusing the person with a film title. This failure stems from constrained local decision-making and an absence of coherent multi-step planning, which leads to retrieval and reasoning errors even when the KG contains the necessary facts. The comparison for previous work and Graph-RFT is in Table 1.
@@ 2. Related Work
SP Methods convert natural language questions into structural queries (e.g., SPARQL) that can be executed on KGs to retrieve answers. Early research efforts (Hu et al., 2017; Xiong et al., 2017; Luo et al., 2025b) in semantic parsing often involved generating query graphs, resembling subgraphs of the KG and allow for direct mapping to logical forms. Lan (Lan and Jiang, 2020) incorporates constraints into the query graph, which effectively narrows the search space and facilitates more flexible query graph generation. Subsequent works (Atif et al., 2023; Schick et al., 2023; Raffel et al., 2020; Wei et al., 2022; Jiang et al., 2022; Lan and Jiang, 2020) utilize sequence-to-sequence models for multi-hop path and SPARQL-expression generation, enhancing the semantic parsing process. Recently, with the emergence of LLMs, various methods unify KGs and LLMs for KGQA.ChatKBQA (Luo et al., 2023a) generates question-aligned logical forms by treating LLMs as semantic parsers and retrieves entities and relations from KGs. RGR-KBQA (Feng and He, 2025) retrieves factual knowledge from KGs to enhance the semantic understanding capabilities of LLMs and finetune them to generate the logical form. Rule-KBQA (Zhang et al., 2025) guides logical form generation via a two-phase process that inducts a rule library and deducts forms via a rule-guided agent. However, these methods depend heavily on the assumption that KGs are complete, which means no answers can be obtained if information is missing.
@@
RA Methods retrieve relevant KG triple based on input questions, then use these triples to guide LLMs to generate accurate answers, emphasizing the dynamic interaction between KG retrieval and LLM reasoning. ToG (Sun et al., 2023) employs an explore-and-exploit approach, which can discover the most promising reasoning paths and return the most likely reasoning results. GoG (Xu et al., 2024) proposes a think-search-generate paradigm, alleviating the incompleteness issue of KGs by using LLMs’ internal knowledge. PoG (Tan et al., 2025b) leverages knowledge reasoning paths as retrieval-augmented inputs for LLMs, thereby alleviating issues related to insufficient domain knowledge and unfaithful reasoning chains. DoG (Ma et al., 2025) introduces a subgraph-focused mechanism that allows LLMs to attempt intermediate answers after each reasoning step, effectively mitigating the drawbacks of lengthy reasoning trajectories. However, most of these approaches depend on powerful closed-source LLM APIs (e.g., GPT-4 (Achiam et al., 2023)), which results in substantial performance degradation when weaker open-source models are used as backbones. KG-Agent (Jiang et al., 2024) develops an iterative reasoning framework that integrates an LLM, a multifunctional toolbox, a KG-based executor, and a knowledge memory module, enabling smaller LLMs to make decisions autonomously. SymAgent (Liu et al., 2025a) conceptualizes the KG as a dynamic environment and is designed to autonomously and effectively integrate the capabilities of both the LLM and the KG. Graph-R1 (Luo et al., 2025a) proposes lightweight knowledge hypergraph construction and retrieval framed as a multi-turn agent–environment interaction, optimizing the reasoning process via an end-to-end reward mechanism. DynaSearcher (Hao et al., 2025) utilizes knowledge graphs as external structured knowledge to guide the search process by explicitly modeling entity relationships, thereby ensuring factual consistency in intermediate queries.
@@ 3.2. CoT Fine-Tuning for Reasoning Activation
where is the question, is the concatenated sequence of reasoning steps and the answer, is the model’s token distribution. The output model serves as the initialization for the next stage (Tan et al., 2025a), ensuring a robust foundation for RL.
@@ 3.3.1. Rollout with Search Tools
Retrieve Facts from the KG. For each single steps obtained through planning, we need to sequentially select relevant triples from the KG. To address this, we design two interactive toolkits.
@@
Given the obtained single-hop questions (e.g., “Identify the country that uses the Iranian Rail”), we first invoke the Relation Search Tool to retrieve a candidate set of relations associated with the initial (topic) entity. The LLM then reasons over this set and selects the most appropriate relation, for example, choosing “location.country.currency_used” to replace “used_in” in the retrieval function. This greedy strategy simplifies the reasoning process by selecting a locally optimal relation at each step, thereby reducing reliance on global contextual information. Next, the Neighbor Search Tool is called to obtain specific triples and their corresponding answers (e.g., “Iran”). If the LLM determines that the retrieved triples are insufficient to answer the question, or if no answer is found in the knowledge graph (i.e., when ¡neighbor_information¿ and ¡/neighbor_information¿ return “No information in the KG, please use web tool”), the system directly invokes the search engine. The answers to higher-priority (according to the order produced by the subquestion plan) single-hop questions can then be combined with those of lower-priority ones to generate new single-hop questions. By iteratively repeating this process, the system is able to derive answers to complex multi-hop questions. Notably, if the maximum iteration limit is reached without successfully generating an answer, the parametric knowledge of the LLM is used to provide a final response.
@@ 3.3.2. Reward Design
| RoG (Luo et al., 2023b) | SL | Qwen2.5-7B-base | 63.7 | 53.6 | 84.2 | 75.3 | - | - | - | - |
| ChatKBQA (Luo et al., 2023a) | SL | Qwen2.5-7B-base | 72.4 | 37.8 | 75.5 | 47.1 | - | - | - | - |
@@ 4. Experiments
Datasets. We evaluate Graph-RFT on four KGQA datasets, including three multi-hop datasets: CWQ (Talmor and Berant, 2018), WebQSP (Yih et al., 2016), GrailQA (Gu et al., 2021) and one single-hop dataset: SimpleQuestions (Petrochuk and Zettlemoyer, ). We consider two scenarios: Complete Knowledge Graph (CKG) and Incomplete Knowledge Graph (IKG). For CKG, we directly use the datasets provided by GoG (Xu et al., 2024). For IKG, since GoG only provides CWQ and WebQSP, we construct IKG versions of GrailQA and SimpleQuestions by sampling 3,000 multi-hop questions from GrailQA and 3,000 questions from SimpleQuestions. We further generate four IKGs with varying completeness levels IKG-20%, IKG-40%, IKG-60%for example, IKG-40% removes 40% of the critical triples for each question and all relations between the corresponding entity pairs. Freebase (Bollacker et al., 2008) serves as the background KG for all datasets.
The impact of using alternative knowledge graphs is discussed in Appendix A.5.
@@ 4.1. Main Results (RQ1)
Table 2 presents our main results on four KGQA datasets. We can observe that Graph-RFT consistently achieves SOTA performance across all settings in IKG, even when built upon smaller backbones such as Qwen2.5-7B-base, outperforming GPT-4-based methods.
@@ 4.3. Different KG Incompleteness (RQ3)
Impact of Rewards with or without Penalty. From Figure 3, Graph-RFT increases the frequency of external searches as KG completeness decreases. This behavior reflects the model’s ability to proactively compensate for missing knowledge required for reasoning through web retrieval. Moreover, the model’s search strategy varies notably with question complexity: compared to WebQSP (maximum 2-hop), Graph-RFT conducts significantly more external searches when tackling the more complex multi-hop questions in CWQ (maximum 4-hop). This demonstrates that the model can dynamically adjust its web search frequency according to question complexity, effectively aligning knowledge acquisition with the reasoning demands of the task.
Impact of Different KG Incompleteness. From Figure 4, we observe that Graph-RFT outperforms GoG across all degrees of incompleteness. This improvement arises because GoG is primarily confined to local decision-making, lacking the global planning needed for effective question understanding and tool invocation in complex problem decomposition. In contrast, when KG triples are missing, GoG relies solely on the LLM’s internal knowledge to generate answers, whereas our model retrieves relevant external documents through web search, thereby supplementing incomplete KG information and improving answer accuracy. Notably, the improvement on CWQ is much more pronounced than on WebQSP, validating Graph-RFT’s strong planning ability and logical decomposition mechanism for tackling complex reasoning tasks.
@@ Acknowledgements.
We thank all anonymous reviewers and area chairs for their comments. This work is supported by the National Natural Science Foundation of China (U23A20316), the CCF-Tencent Rhino-Bird Open Research Fund (CCF-Tencent RAGR20250115), and the computing power resources from Wuhan Dongxihu District Intelligent Computing Center.References
- Gpt-4 technical report. arXiv preprint arXiv:2303.08774. Cited by: §2.
@@
- Beamqa: multi-hop knowledge graph question answering with sequence-to-sequence prediction and beam search. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, pp. 781–790. Cited by: §2.
@@
- Knowledge-augmented language model prompting for zero-shot knowledge graph question answering. arXiv preprint arXiv:2306.04136. Cited by: §4.
@@
- Freebase: a collaboratively created graph database for structuring human knowledge. In Proceedings of the 2008 ACM SIGMOD international conference on Management of data, pp. 1247–1250. Cited by: §4.
- Think you have solved question answering? try arc, the ai2 reasoning challenge. arXiv preprint arXiv:1803.05457. Cited by: §1.
- Discourse on method. Aladdin Book Company. Cited by: §1.
@@
- Rgr-kbqa: generating logical forms for question answering using knowledge-graph-enhanced large language model. In Proceedings of the 31st International Conference on Computational Linguistics, pp. 3057–3070. Cited by: Table 1, §2.
@@
- Beyond iid: three levels of generalization for question answering on knowledge bases. In Proceedings of the web conference 2021, pp. 3477–3488. Cited by: §4.
- Deepseek-r1: incentivizing reasoning capability in llms via reinforcement learning. arXiv preprint arXiv:2501.12948. Cited by: §1, §1.
@@
- DynaSearcher: dynamic knowledge graph augmented search agent via multi-reward reinforcement learning. arXiv preprint arXiv:2507.17365. Cited by: §2, §3.3.2.
- Measuring massive multitask language understanding. arXiv preprint arXiv:2009.03300. Cited by: §1.
@@
- Answering natural language questions by subgraph matching over knowledge graphs. IEEE Transactions on Knowledge and Data Engineering 30 (5), pp. 824–837. Cited by: §2.
@@
- Structgpt: a general framework for large language model to reason over structured data. arXiv preprint arXiv:2305.09645. Cited by: §A.2, §1, Table 2, §4.
@@
- Kg-agent: an efficient autonomous agent framework for complex reasoning over knowledge graph. arXiv preprint arXiv:2402.11163. Cited by: Table 1, §2.
@@
- Unikgqa: unified retrieval and reasoning for solving multi-hop question answering over knowledge graph. arXiv preprint arXiv:2212.00959. Cited by: §2.
@@
- Search-r1: training llms to reason and leverage search engines with reinforcement learning. arXiv preprint arXiv:2503.09516. Cited by: §3.3.2.
- Research trends for the interplay between large language models and knowledge graphs. arXiv preprint arXiv:2406.08223. Cited by: §1.
@@
- Query graph generation for answering multi-hop complex questions from knowledge bases. Cited by: §2.
@@
- Few-shot in-context learning for knowledge base question answering. arXiv preprint arXiv:2305.01750. Cited by: §A.2, §1, Table 2.
@@
- Chain of knowledge: a framework for grounding large language models with structured knowledge bases. arXiv preprint arXiv:2305.13269 3. Cited by: §4.
@@
- Symagent: a neural-symbolic self-learning agent framework for complex reasoning over knowledge graphs. In Proceedings of the ACM on Web Conference 2025, pp. 98–108. Cited by: §1, §2.
- Visual-rft: visual reinforcement fine-tuning. arXiv preprint arXiv:2503.01785. Cited by: §1.
- Graph-r1: towards agentic graphrag framework via end-to-end reinforcement learning. arXiv preprint arXiv:2507.21892. Cited by: §2.
@@
- Chatkbqa: a generate-then-retrieve framework for knowledge base question answering with fine-tuned large language models. arXiv preprint arXiv:2310.08975. Cited by: §A.2, §1, §2, Table 2.
@@
- Chatrule: mining logical rules with large language models for knowledge graph reasoning. In Pacific-Asia Conference on Knowledge Discovery and Data Mining, pp. 314–325. Cited by: §2.
@@
- Reasoning on graphs: faithful and interpretable large language model reasoning. arXiv preprint arXiv:2310.01061. Cited by: §A.2, Table 2.
@@
- Debate on graph: a flexible and reliable reasoning framework for large language models. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 39, pp. 24768–24776. Cited by: Table 1, §1, §2.
@@
- [28] Simplequestions nearly solved: a new upperbound and baseline approach. arxiv 2018. arXiv preprint arXiv:1804.08798. Cited by: §4.
@@
- Exploring the limits of transfer learning with a unified text-to-text transformer. Journal of machine learning research 21 (140), pp. 1–67. Cited by: §2.
- Toolformer: language models can teach themselves to use tools. Advances in Neural Information Processing Systems 36, pp. 68539–68551. Cited by: §2.